# SymmCD: Symmetry-Preserving Crystal Generation with Diffusion Models

### v1

@@ 5 Experiments
We empirically demonstrate our contributions, particularly in ensuring we generate crystals with desired symmetries while being competitive with existing baselines. In other words, we show that SymmCD generates symmetric, stable, and valid crystals. We compare our proposed method with four recent strong baselines: CDVAE (Xie et al., 2022), DiffCSP (Jiao et al., 2023), DiffCSP++ (Jiao et al., 2024) and FlowMM (Miller et al., 2024).
@@ 5.1 Symmetry and structural diversity
Table 1: Template statistics for various models.
@@
To detect the space group of the generated structures, we use spglib’s symmetry finding method (Togo & Tanaka, 2018a; Ong et al., 2013) with a tolerance of 0.1Å. This is applied to 10,000 crystals sampled from each model. The distribution of space groups of the generated structures is shown in Figure 4. It can be observed that while SymmCD matches the highly diverse data distribution, CDVAE mostly generates crystals with trivial $P1$ symmetry, and DiffCSP and FlowMM generate many crystals with low symmetry and generally have lower diversity of space groups. We also consider a new quantitative metric to characterize the space group distribution, $d_{\mathrm{sg}}$, which is calculated as the Jensen-Shannon distance between the distribution of space groups of the generated structures and the test set. We report it for the different methods in the rightmost column of Table 2. The results confirm that SymmCD and DiffCSP++ are the only methods that accurately match the distribution of space groups in the dataset.
@@
The training dataset contains 3318 such unique templates. We examine the templates for the 10,000 crystals generated by each method, and report results in Table 1. We find that SymmCD performs best out of all models, proposing unique and novel templates. This highlights an important limitation of DiffCSP++. While it is able to produce diverse space groups and to a certain extent diverse templates, since it uses pre-defined templates it fails to generate structures with novel templates. Our method does not suffer from this problem since it learns to generate templates.
5.2 Proxy metrics
We compare the different methods using the metrics established by Xie et al. (2022), measuring the validity, coverage, and property statistics of the generated crystals. We measure the validity by checking structural validity, defined as whether no two atoms are closer than 0.5 Å apart, and compositional validity, defined as whether the charges are balanced as determined by SMACT (Davies et al., 2019) [^3] It should be noted that the compositional validity of the MP-20 dataset is only 92%.. To determine coverage, we examine the CrystalNN structural fingerprints (Zimmermann & Jain, 2020) and Magpie compositional fingerprints (Ward et al., 2016) of the generated crystals, and look at their distances to the fingerprints of the crystals in the test set. This gives us recall and precision metrics. We look at the distances between the properties of the generated crystals and the crystals from the test set to compare the ability of each model to match the data distribution. We specifically compare the Wasserstein distances between the atomic densities $d_{\rho}$, number of unique elements $d_{\mathrm{elem}}$, and predicted formation energy $d_{\mathrm{E}}$. The results are shown in Table 2. We observe that SymmCD performs on par with other methods across different metrics, and that sampling from a smaller set of space groups improves the validity of crystals while trading off diversity and matching the data distribution. These results also show that SymmCD can generalize to generate valid structures even for groups which are rarely represented in the training data. We also consider a variant of the model (10 SGs) where we sample space groups from the MP-20 distribution, restricted to the 10 most common space groups, similar to (Cao et al., 2024)  [^4] These space groups are numbered: 2, 12, 14, 62, 63, 139, 166, 194, 221, 225 . This is to provide a more nuanced comparison with other methods, which are not constrained in matching the space group distribution. This choice still captures a large portion of the data distribution, since these are the most prevalent space groups.
Table 2: Results for comparing the validity, coverage, and property distribution metrics. Best results in each category are bolded.
Struct. Comp. Recall Precision $d_{\rho}$ $d_{\mathrm{E}}$ $d_{\mathrm{elem}}$ $d_{\mathrm{sg}}$
CDVAE 99.93 86.93 98.31 99.35 0.9144 0.1645 1.6538 0.7263
DiffCSP 99.61 82.23 99.53 99.35 0.2565 0.1402 0.4027 0.4446
DiffCSP++ 99.99 85.81 99.48 99.66 0.2779 0.0872 0.4079 0.0771
FlowMM 96.43 83.37 99.47 99.71 0.2905 0.1072 0.0788 0.5137
SymmCD (All SGs) 94.32 85.85 99.64 98.87 0.0901 0.1166 0.3990 0.0899
SymmCD (10 SGs) 97.31 87.10 97.21 99.42 0.2829 0.1510 0.1769 0.4737
Table 3: Number of stable and S.U.N. samples produced from an initial set of 1000 generated crystals for each method.
@@ 5.2 Proxy metrics
CDVAE 0.1% 3.6% 3.5%
DiffCSP 8.9% 12.5% 9.7%
DiffCSP++ 8.9% 13.2% 9.1%
FlowMM 4.1% 9.3% 6.3%
SymmCD (all SGs) 5.0% 9.4% 7.0%
SymmCD (10 SGs) 7.9% 11.7% 9.9%
@@
Regardless of their target application, generative models for crystals should produce sets of crystals that are thermodynamically stable, unique (not duplicated within the predicted set), and novel (not already in the training data), or S.U.N. To this end, we adapt the evaluation procedure of Miller et al. (2024) to assess the capability of our model to generate S.U.N. materials. Thermodynamic stability is determined by estimating the energy of a material with respect to a convex hull. The convex hull gives linear combinations of known phases that represent the lowest-energy mixtures of materials; if a material has an energy above the hull, it is energetically favorable for it decompose into a combination of these stable phases and is therefore thermodynamically unstable. We assess the stability of generated crystals by estimating their energies using a pretrained CHGNet model (Deng et al., 2023), and comparing that to a convex hull computed for Materials Project (Riebesell et al., 2024).
@@
For each method, we randomly sub-sample 1000 crystals of the 10,000 generated samples and predict their stability. We also use CHGNet to compute relaxed structures for each crystal, which results in higher stability. Finally, we check whether the stable relaxed crystals are also unique and novel. Details of this procedure are included in Miller et al. (2024). Note that we use a machine learning potential instead of a full Density Functional Theory (DFT) calculation, as DFT relaxation would be orders of magnitude more expensive to compute.
The results are shown in Table 3. SymmCD (all SGs) performs slightly better than FlowMM in generating stable structures, but worse than DiffCSP and DiffCSP++. The version sampling from a smaller number of space groups however obtains a larger proportion of S.U.N. structures than all baselines. Note that, while DiffCSP++ has a larger proportion of relaxed stable structures, filtering for unique and novel structures gives SymmCD the advantage, providing more evidence that it generates more diverse structures.
@@ 5.4 Computational efficiency
Table 4: Computational efficiency of our compact representation with a 40 GB NVIDIA MIG A100 instance.
@@
Time for one training epoch ($\downarrow$) 27 sec. 52 sec.
Finally, we demonstrate significant computational efficiency gains and reduced memory footprint due to using a more compact representation based on crystallographic orbits. We compare our model to an equivalent model that looks at a full unit cell, rather than just the asymmetric unit. It also uses a fully connected graph to represent the atoms in the unit cell, but unlike SymmCD, it does not use site symmetry representations as they are not necessary. This makes the model essentially equivalent to DiffCSP (Jiao et al., 2023), but with the same architecture and hyperparameters as SymmCD for consistent comparison. We compare the two representations for an epoch of training using 40GB of RAM and a single NVIDIA MIG A100 instance and report the results in Table 4. These results highlight SymmCD’s memory efficiency and faster training capabilities.
@@ Conclusion
In this paper, we introduced a novel approach for generating crystals with precise symmetry properties. We proposed to leverage asymmetric units and site-symmetry representations within a diffusion model framework. This approach ensures that the generated crystals inherently preserve desired symmetries while allowing greater diversity, computational efficiency and flexibility in the generation process. To encode crystal and site symmetries we introduced a new representation of crystal symmetries that enables information sharing across space groups, improving generalization when learning with a diverse set of crystal symmetries. Our results indicate that this method produces valid, stable, novel, and structurally diverse crystals, and improves computational efficiency of the generative model showing promise for discovery in materials science. In this work, we focused on inorganic crystals, but SymmCD could potentially be promising for applications on molecular crystals and co-crystals, which also have non-trivial symmetries. Beyond materials, other data types such as molecules and graphs often exhibit complex symmetries. Future work could investigate if symmetry constraints could also be useful in these modalities to improve generative models.
One limitation of our framework is that it makes it more challenging to perform crystal structure prediction given a composition, since it relies on sampling a space group first, and then a composition conditioned on the space group. Finally, an important area of future work in generative models for crystals is also to go beyond single crystals, and consider generation of polycrystalline materials. These types of materials are common in applications, yet not suited to generation using single unit cells or asymmetric units.
@@ Appendix D Diffusion and denoising process details
We perform diffusion over the fractional coordinates using the same method as (Jiao et al., 2023). Due to the periodicity of fractional coordinates, the noising process $q({\mathbf{X}}_{t}|{\mathbf{X}}_{0})$ is determined by a Wrapped Normal distribution rather than a Gaussian distribution, and we initialize the fractional coordinates ${\mathbf{X}}_{T}$ with the uniform distribution $U(0,1)$ when sampling.
@@
We use discrete diffusion from Austin et al. (2021) to sample the atom types of each representative. If $\mathbf{a}_{0}\in\{0,1\}^{Z}$ is the one-hot encoding of atom types for a single representative, then we can noise it as: $q(\mathbf{a}_{t}|\mathbf{a}_{0})=\mathrm{Cat}(\mathbf{a}_{t};\mathbf{p}=\mathbf{a}_{0}^{\top}\bar{\mathbf{Q}}_{t})$, where $\bar{\mathbf{Q}}_{t}=\prod_{i=1}^{t}\mathbf{Q}_{i}\in\mathbb{R}^{Z\times Z}$ is the cumulative product of transition matrices between timesteps. Inspired by Vignac et al. (2023), the transition matrix can be parametrized quite simply as $\mathbf{Q}_{t}=\alpha_{t}\mathbf{I}+\beta_{t}\mathbf{m}_{a}$, where $\mathbf{m}_{a}$ are the marginals over the atom types in the data, and $\alpha_{t}$ and $\beta_{t}$ are scheduling parameters. The effect of this noising scheme is that regardless of $\mathbf{a}_{0}$, the fully noised $\mathbf{a}_{T}=\mathbf{a}_{0}^{\top}\mathbf{Q}_{T}=\mathbf{m}_{a}$, so we can sample from the prior distribution $\mathbf{m}_{a}$, which is close to the data distribution. The discrete diffusion model is trained using a cross-entropy loss:
@@
Diffusion for site-symmetries ${\mathbf{S}}$
The site-symmetry representation matrices described in Section 4.2 can be thought of as 15 separate 13-dimensional categorical variables: one site-symmetry operation per axis. Our diffusion model over site-symmetries is almost identical to the method for atom types, applying discrete diffusion separately over each of the axes. Because the site-symmetries depend strongly on the space group, we use transition matrices that are different for each space group: $\mathbf{Q}_{t,i,G}=\alpha_{t}\mathbf{I}+\beta_{t}\mathbf{m}_{S_{\alpha},G}$, where $\mathbf{m}_{S_{u},G}$ denotes the marginals over site-symmetry operations for axis $S_{u}$ given space group $G$. For each representative node, we average the cross-entropy loss over each of the axes.
@@ E.2 Model Hyperparameters
Figure 5: Histogram of number of atoms in crystals from MP-20 and generated by SymmCD.
Appendix F Additional Results
F.1 Number of atoms
To demonstrate that SymmCD is able to correctly predict reasonable site symmetries, we show here that the distribution of number of atoms per crystal matches the dataset it is trained on. This is not a trivial task, as the model needs to learn the multiplicity of different possible site symmetries, which depends on both the different symmetry elements of the site symmetry and the space group that it belongs to.
F.2 Property Prediction task

### v2

@@ 5 Experiments
We empirically demonstrate our contributions, particularly in ensuring we generate crystals with desired symmetries while being competitive with existing baselines. We compare our proposed method with four recent strong baselines: CDVAE (Xie et al., 2022), DiffCSP (Jiao et al., 2023), DiffCSP++ (Jiao et al., 2024) and FlowMM (Miller et al., 2024). We retrained each method according to their given hyperparameters, and generated 10,000 crystals each. We also consider a variant of our model (10 SGs) where we only sample from the 10 most common space groups in the MP-20 dataset, similar to (Cao et al., 2024)  [^3] These space groups are numbered: 2, 12, 14, 62, 63, 139, 166, 194, 221, 225 . This is to provide a more nuanced comparison with other methods, which are not constrained in matching the space group distribution. This choice still captures a large portion of the data distribution, since these are the most prevalent space groups.
@@ 5.1 Symmetry and structural diversity
Table 1: Template statistics for each model.
@@
To detect the space group of the generated structures, we use spglib’s symmetry finding method (Togo & Tanaka, 2018a; Ong et al., 2013) with a tolerance of 0.1Å. The distribution of space groups of the generated structures is shown in Figure 4. It can be observed that while SymmCD matches the highly diverse data distribution, CDVAE mostly generates crystals with trivial $P1$ symmetry, and DiffCSP and FlowMM generate many crystals with low symmetry and generally have lower diversity of space groups. We also consider a new quantitative metric to characterize the space group distribution, $d_{\mathrm{sg}}$, which is calculated as the Jensen-Shannon distance between the distribution of space groups of the generated structures and the test set. We report it for the different methods in the rightmost column of Table 3. The results confirm that SymmCD and DiffCSP++ are the only methods that accurately match the distribution of space groups in the dataset.
@@
The training dataset contains 3318 such unique templates. We examine the templates for the 10,000 crystals generated by each method, and report results in Table 1. We find that SymmCD performs best out of all models, proposing unique and novel templates. This highlights an important limitation of DiffCSP++. While it can produce diverse space groups and to a certain extent diverse templates, since it uses pre-defined templates it fails to generate structures with novel templates. Our method does not suffer from this problem since it learns to generate templates.
Table 2: Percent of stable and S.U.N. samples produced from an initial set of 10,000 generated crystals for each method.
@@
CDVAE 0.24% 4.42% 4.26%
DiffCSP 7.82% 11.32% 8.92%
DiffCSP++ 6.99% 11.36% 8.62%
FlowMM 4.26% 9.05% 6.49%
SymmCD (All SGs) 4.96% 9.34% 6.86%
SymmCD (10 SGs) 6.85% 10.92% 7.59%
@@
Regardless of their target application, generative models for crystals should produce sets of crystals that are thermodynamically stable, unique (not duplicated within the predicted set), and novel (not already in the training data), or S.U.N. To this end, we adapt the evaluation procedure of Miller et al. (2024) to assess the capability of our model to generate S.U.N. materials. Thermodynamic stability is determined by estimating the energy of a material with respect to a convex hull. The convex hull gives linear combinations of known phases that represent the lowest-energy mixtures of materials; if a material has an energy above the hull, it is energetically favorable for it to decompose into a combination of stable phases and is therefore thermodynamically unstable. We assess the stability of generated crystals by estimating their energies using a pretrained CHGNet model (Deng et al., 2023), and comparing that to a convex hull computed for Materials Project (Riebesell et al., 2024).
@@
We predict the stability of the 10,000 samples generated by each method, and then use CHGNet to compute relaxed structures for each crystal, which results in higher stability. Finally, we check whether the stable relaxed crystals are also unique and novel. Details of this procedure are included in Miller et al. (2024). The results are shown in Table 2. We see that SymmCD performs on par with the baselines. Sampling over the most common space groups results in more stable structures.
5.3 Validation with density functional theory relaxations
We carried out electronic structure calculations as a more accurate way to evaluate the crystal structures generated by the different methods. Since these computations are significantly more expensive than the ones with CHGNet, we performed them on 100 structures sampled from each model. Concretely, these involved using the CP2K (Kühne et al., 2020) suite of programs to perform cell and geometry relaxations to assess the quality of generated structures based on their distances from “true” local minima. We expect a better method to produce structures with less distance traveled, smaller per-atom forces (i.e. energy gradient with respect to positions), and with fewer iterations.
In these calculations, we find that the convergence rates of methods are not significantly different (see Figure 10). Comparing expectation values of the maximum force after optimization, however, we see that SymmCD is significantly more successful in generating structures that are readily optimized. The full results of this analysis, as well as the parameters used for CP2K can be found in Section F.6.
5.4 Proxy metrics
Although ultimately we care about the stability, novelty, and properties of crystals after structural relaxation, we also compare the different methods using the proxy metrics established by Xie et al. (2022), as they are cheap enough to compute over a large set of generated samples and demonstrate the ability of different methods to capture the properties of a target distribution. These metrics include heuristics of validity, coverage of the test set, and the Wasserstein distances between the distributions of three properties of the generated samples and the test set: atomic densities $d_{\rho}$, number of unique elements $d_{\mathrm{elem}}$, and predicted formation energy $d_{\mathrm{E}}$. We also include $d_{\mathrm{sg}}$, the Jensen-Shannon distance between space group distributions. Further details are included in Section F.5.
For each method, we retrained a model with 5 different seeds, generating 10,000 samples for evaluation per seed. The results are shown in Table 3. Although SymmCD performs similarly to other methods for most metrics, it is more likely to generate structurally invalid crystals, possibly because it cannot easily see distances between atoms. We also note the large variance in the results across different seeds for each method.
Table 3: Results for comparing the validity, coverage, and property distribution metrics.
MP-20 Struct. Comp. Recall Precision $d_{\rho}$ $d_{E}$ $d_{\mathrm{elem}}$ $d_{\mathrm{sg}}$
CDVAE $99.97_{\pm 0.03}$ $85.61_{\pm 1.20}$ $99.31_{\pm.06}$ $99.47_{\pm 0.19}$ $0.70_{\pm.10}$ $0.24_{\pm.06}$ $1.28_{\pm.06}$ $0.69_{\pm.007}$
DiffCSP $97.43_{\pm 3.10}$ $82.50_{\pm 1.34}$ $99.55_{\pm.08}$ $98.73_{\pm 1.34}$ $0.18_{\pm.08}$ $0.14_{\pm.05}$ $0.56_{\pm.07}$ $0.44_{\pm.009}$
DiffCSP++ $99.44_{\pm 0.06}$ $86.50_{\pm 0.85}$ $99.72_{\pm.06}$ $99.61_{\pm 0.08}$ $0.12_{\pm.04}$ $0.05_{\pm.01}$ $0.33_{\pm.04}$ $0.16_{\pm.009}$
FlowMM $96.67_{\pm 0.57}$ $83.25_{\pm 0.13}$ $99.49_{\pm.05}$ $99.58_{\pm 0.10}$ $0.23_{\pm.10}$ $0.09_{\pm.02}$ $0.08_{\pm.02}$ $0.50_{\pm.011}$
SymmCDAll ${}_{\text{All SGs}}$ $90.34_{\pm 3.39}$ $85.81_{\pm 0.87}$ $99.58_{\pm.08}$ $97.76_{\pm 0.85}$ $0.23_{\pm.06}$ $0.21_{\pm.04}$ $0.40_{\pm.06}$ $0.164_{\pm.003}$
SymmCD10 ${}_{\text{10 SGs}}$ $92.30_{\pm 6.10}$ $87.13_{\pm 0.62}$ $97.33_{\pm.59}$ $98.78_{\pm 0.69}$ $0.53_{\pm.23}$ $0.21_{\pm.09}$ $0.16_{\pm.02}$ $0.469_{\pm.002}$
@@ 5.5 Computational efficiency
Table 4: Computational efficiency of our representation
@@
Time for 1 training epoch ($\downarrow$) 27 sec. 52 sec.
We demonstrate significant computational efficiency gains and reduced memory footprint due to using a more compact representation based on crystallographic orbits. We compare our model to an equivalent model that looks at a full unit cell, rather than just the asymmetric unit. It also uses a fully connected graph to represent the atoms in the unit cell, but does not predict site symmetry representations. This makes the model essentially equivalent to DiffCSP (Jiao et al., 2023), but with the same architecture and hyperparameters as SymmCD for consistent comparison. We compare the two representations for one epoch of training using 40GB of RAM and a single NVIDIA MIG A100 and report the results in Table 4. These results highlight SymmCD’s memory efficiency and faster training capabilities.
5.6 Scaling up to MPTS-52
To validate that SymmCD can scale up to datasets of larger crystals thanks to its compact representation, we also validate our model by training it on the more challenging MPTS-52 dataset (Baird et al., 2024), which contains crystals with up to 52 atoms per primitive unit cell, and includes more chemical elements. Further details of this dataset along with the results are included in Section F.1.
We note that none of the baselines we compare against (CDVAE, DiffCSP, FlowMM, and DiffCSP++) present results for de-novo generation on this dataset. We can see from these results that SymmCD is able to scale to much larger crystals than those in the MP-20 dataset and is still capable of producing valid, stable, and novel crystals that match the training distribution of the dataset. The metrics reported are worse than those for MP-20, as this is a more difficult dataset.
@@ Conclusion
In this paper, we introduced a novel approach for generating crystals with precise symmetry properties. We proposed to leverage asymmetric units and site symmetry representations within a diffusion model framework. This approach ensures that the generated crystals inherently preserve desired symmetries while allowing greater diversity, computational efficiency, and flexibility in the generation process. To encode crystal and site symmetries we introduced a new representation of crystal symmetries that enables information sharing across space groups, improving generalization when learning with a diverse set of crystal symmetries. Our results indicate that this method produces stable, novel, and structurally diverse crystals, with improved computational efficiency, showing promise for discovery in materials science. In this work, we focused on inorganic crystals, but SymmCD could potentially be promising for applications on molecular crystals and co-crystals, which also have non-trivial symmetries. Beyond materials, other data types such as molecules and graphs often exhibit complex symmetries, and future work could investigate if symmetry constraints could also be useful in generative models for these modalities.
A limitation of our framework is that it makes it more challenging to perform crystal structure prediction given a composition since it relies on sampling a space group first, and then a composition conditioned on the space group. Finally, an important area of future work in generative models for crystals is also to go beyond single crystals, and consider the generation of polycrystalline materials. These types of materials are common in applications, yet not suited to generation using single unit cells or asymmetric units.
@@ Appendix D Diffusion and denoising process details
We perform diffusion over the fractional coordinates using the same method as Jiao et al. (2023). Due to the periodicity of fractional coordinates, the noising process $q({\mathbf{X}}_{t}|{\mathbf{X}}_{0})$ is determined by a Wrapped Normal distribution rather than a Gaussian distribution, and we initialize the fractional coordinates ${\mathbf{X}}_{T}$ with the uniform distribution $U(0,1)$ when sampling.
@@
We use discrete diffusion from Austin et al. (2021) to sample the atom types of each representative. If $\mathbf{a}_{0}\in\{0,1\}^{Z}$ is the one-hot encoding of atom types for a single representative, then we can noise it as: $q(\mathbf{a}_{t}|\mathbf{a}_{0})=\mathrm{Cat}(\mathbf{a}_{t};\mathbf{p}=\mathbf{a}_{0}^{\top}\bar{\mathbf{Q}}_{t})$, where $\bar{\mathbf{Q}}_{t}=\prod_{i=1}^{t}\mathbf{Q}_{i}\in\mathbb{R}^{Z\times Z}$ is the cumulative product of transition matrices between timesteps. Inspired by Vignac et al. (2023), the transition matrix can be parametrized as $\mathbf{Q}_{t}=\alpha_{t}\mathbf{I}+\beta_{t}\mathbf{m}_{a}$, where $\mathbf{m}_{a}$ are the marginals over the atom types in the data, and $\alpha_{t}$ and $\beta_{t}$ are scheduling parameters. The effect of this noising scheme is that regardless of $\mathbf{a}_{0}$, the fully noised $\mathbf{a}_{T}=\mathbf{a}_{0}^{\top}\mathbf{Q}_{T}=\mathbf{m}_{a}$, so we can sample from the prior distribution $\mathbf{m}_{a}$, which is close to the data distribution. The discrete diffusion model is trained using a cross-entropy loss:
@@
Diffusion for site symmetries ${\mathbf{S}}$
The site symmetry representation matrices described in Section 4.2 can be thought of as 15 separate 13-dimensional categorical variables: one site symmetry operation per axis. Our diffusion model over site symmetries is almost identical to the method for atom types, applying discrete diffusion separately over each of the axes. Because the site symmetries depend strongly on the space group, we use transition matrices that are different for each space group: $\mathbf{Q}_{t,i,G}=\alpha_{t}\mathbf{I}+\beta_{t}\mathbf{m}_{S_{\alpha},G}$, where $\mathbf{m}_{S_{u},G}$ denotes the marginals over site symmetry operations for axis $S_{u}$ given space group $G$. For each representative node, we average the cross-entropy loss over each of the axes.
@@ E.2 Model Hyperparameters
Appendix F Extended Results
F.1 MPTS 52 Dataset
[…]
In addition to MP-20, we also trained SymmCD on the MPTS-52 dataset (Baird et al., 2024), a more challenging subset of the Materials Project that contains materials with up to 52 atoms per primitive unit cell. Unlike MP-20, it does not filter out materials containing radioactive elements. The dataset contains 40,476 samples with a train/validation/test split of 27,380/5,000/8,096 crystals. The splits are in chronological order, with the materials in the test set having been discovered most recently, and the materials in the training set having been discovered earliest. The distribution of space groups found in the dataset are shown in Figure 5. None of the diffusion and flow-based methods we compared against (CDVAE, DiffCSP, DiffCSP++, or FlowMM) have reported de-novo generation results on this dataset.
We trained SymmCD on MPTS-52 using all of the same hyperparameters as were used for the MP-20 dataset, but trained for 1500 epochs. We sampled 10,000 crystals, and checked the same proxy metrics. We also relaxed the generated crystals using CHGNet and checked for whether the generated crystals were S.U.N. The unique templates (as described in Section 5.1) of the dataset and of our generated crystals are shown in Table 7. The results for the proxy metrics are shown in Table 7 and stability results are shown in LABEL:tab:mpts52_sun.
Table 5: Template statistics for MPTS-52
Method # Unique % in Train # New
Training Set 4452 100 -
SymmCD 2772 48.3 1432
Table 6: Percent of stable and S.U.N. samples produced from an initial set of 10,000 generated crystals for SymmCD trained on the MPTS-52 dataset.
Initial Relaxed Relaxed
Stable Stable S.U.N.
1.72% 5.97% 4.62%
Table 7: The validity, coverage, and property distribution metrics for SymmCD trained on the MPTS-52 dataset.
Struct. Comp. Recall Precision $d_{\rho}$ $d_{E}$ $d_{\mathrm{elem}}$ $d_{\mathrm{sg}}$
$90.1$ $79.2$ $99.6$ $96.3$ $0.844$ $0.549$ $0.317$ $0.274$
F.2 Number of atoms
Figure 6: Histogram of the number of atoms in crystals from MP-20 and generated by SymmCD when trained on MP-20.
To demonstrate that SymmCD is able to correctly predict reasonable site symmetries, we show here that the distribution of number of atoms per crystal matches the dataset it is trained on. This is not a trivial task, as the model needs to learn the multiplicity of different possible site symmetries, which depends on both the different symmetry elements of the site symmetry and the space group that it belongs to. The comparison for MP-20 is shown in Figure 6, and the comparison for MPTS-52 is shown in Figure 7.
Figure 7: Histogram of the number of atoms in crystals from MPTS-52 and generated by SymmCD when trained on MPTS-52.
F.3 Property Prediction task